OpenStudy (anonymous):

Expand (multiply) the following polynomial: (x + 3)2(x + 4) + (x + 3)3

OpenStudy (anonymous):

\[ (x + 3)^2(x + 4) + (x + 3)^3\]

OpenStudy (jack1):

k use foil on the first 2, then add the resulting answer and the final brackets...?

OpenStudy (jack1):

(x + 3)2(x + 4) (x + 3)(x + 3)(x + 4) (x + 3)(x + 3) = x^2 + 3x + 3x + 9 (x^2 + 6x + 9)(x + 4) x^3 + 6x^2 + 9x + 4x^2 + 24x + 36 add to final bracket

OpenStudy (anonymous):

but where does the (x+3)^3 go?

OpenStudy (jack1):

(x + 3)3 expand this out using foil... add to other brackets <x^3 + 6x^2 + 9x + 4x^2 + 24x + 36>

OpenStudy (anonymous):

yeah but the solution here...

OpenStudy (jack1):

where?

OpenStudy (anonymous):

(x+ 3)2(x + 4) + (x + 3)3 = (x + 3)2[(x + 4) + (x + 3)] = (x + 3)2[x + 4 + x + 3] = x2 + 6x + 9(2x + 7) = x2(2x + 7) + 6x(2x + 7) + 9(2x + 7) = x2(2x) + x2(7) + 6x(2x) + 6x(7) + 9(2x) + 9(7) = 2x3 + 7x2 + 12x2 + 42x + 18x + 63 = 2x3 + 19x2 + 60x + 63

OpenStudy (anonymous):

I don't understand this part (x+ 3)^2(x + 4) + (x + 3)^3 = (x + 3)^2[(x + 4) + (x + 3)] = where does the root 3 go?

OpenStudy (jack1):

(x + 3)2(x + 4) (x + 3)(x + 3)(x + 4) (x + 3)(x + 3) = x^2 + 3x + 3x + 9 so (x^2 + 6x + 9)(x + 4) x^3 + 6x^2 + 9x + 4x^2 + 24x + 36 now (x+3)^3 (x+3)^3 = (x+3)(x+3)(x+3) = (x^2 + 3x + 3x + 9)(x+3) = (x^3 + 6x^2 + 9x + 3x^2 + 18x + 27 final answer = x^3 + 6x^2 + 9x + 4x^2 + 24x + 36 + x^3 + 6x^2 + 9x + 3x^2 + 18x + 27 = 2x^3 + 12x^2 + 18x + 7x^2 + 42x + 63 = 2x^3 + 19x^2 + 60x + 63 ... i think...?

OpenStudy (jack1):

yeah that doesnt make sense to me about the rt 3, sorry ???

OpenStudy (anonymous):

yeah... it says the answer is 2x^3 + 19x^2 + 60x + 63

OpenStudy (anonymous):

so yeah idk what their talking about with the root 3 but you got the answer the way I calculated mine so I guess it doesn't matter...

OpenStudy (jack1):

sweet, coz im totes lost h=there hey , sorry

OpenStudy (anonymous):

No its cool, thanks for the help.